A body is freely falling from a certain height. In the last second it covers a distance 7h/16 where h is height max. Find H max

Grade:12th pass

1 Answers

Vikas TU
7 years ago
Total distance say in t seconds,
H = 0 + 0.5gt^2
and 
in t-1 seonds,
h = 0 + 0.5g(t-1)^2
Now given,
H -h = 7h/16 = 0.5gt^2 – 0.5g(t-1)^2
solving we get 
t = 7h/16 + ½
put t in H we get,
H = 0.5g(7h/16 + ½)^2 meter.